how many moles of propane, c3h8 is burned to produce 2.00 moles of co2 gas?
1
Expert's answer
2020-11-26T05:09:40-0500

C3H8+5O2=3CO2+4H2O

n (C3H8)=1/3 x n (CO2) = 1/3 x 2=0.67 mol